I am confused about two parts of this equation:
I am given one function and I am supposed to find the other five functions with that given function.
The function I was given is:
cos A = 1/3
I have already found the following functions:
sin A = 2 sqrt(2)/3
tan A = 2 sqrt(2)
sec A = 3
My answers for functions "cot A" and "csc A" are:
cot A = 1/ 2 sqrt(2)
csc A = 3/ 2 sqrt(2)
The answers for "cot A" and "csc A" on the answer sheet are:
cot A = sqrt(2)/4
csc A = 3 sqrt(2)/4
I am confused about how the answers were found.

Answer by lwsshak3(11628)

Your answers are correct except you need to rationalize the denominators to make them agree with the book answers.
cot A = 1/ 2 sqrt(2)*√2/√2=√2/4
csc A = 3/ 2 sqrt(2)*√2/√2=3√2/4